# A Survey on Point-of-Interest Recommendation: Models, Architectures, and Security

### v1

@@ I Introduction
The proliferation of smart devices has fueled the rapid growth of Location-based Social Networks (LBSNs) [1, 2, 3], which allow users to share check-ins, reviews, and personal experiences tied to specific locations. These networks, now with billions of users, generate vast amounts of spatio-temporal data [4, 5, 6, 7, 8, 9], presenting valuable opportunities for personalized Point-of-Interest (POI) recommendations. As a dynamic area in recommendation systems, POI recommendation has gained considerable interest from both users and businesses in recent years. These methods leverage users’ historical check-ins along with multimodal data to suggest personalized destinations [6, 10]. However, the diversity in data size, modality, and user expectations introduces new challenges. These complexities motivate researchers to develop innovative techniques that effectively capture mobility patterns and other relevant features, such as spatial, social, and textual information, to enhance the performance of POI recommendations [11, 12, 13].
@@
While existing surveys [34, 35, 36, 37] have contributed valuable insights into POI recommendation systems, there remains a critical need for a comprehensive review that reflects the rapid developments across POI models, architectures, and security. For instance, while Zhao et al. [37] provide a thorough review of POI recommendation with traditional techniques like matrix factorization, the growing challenges within deep POI recommender systems are not covered. By contrast, Wang et al. [34] provide an overview of various POI recommendation methods in the deep learning era, yet they do not delve deeply into the architectural and security challenges that have emerged with the rise of decentralized systems and practical privacy concerns. Similarly, Islam’s survey [35] emphasizes the impact of deep learning on POI recommendation, but it overlooks key advancements in federated learning and edge computing, which are increasingly shaping the deployment of these systems. Furthermore, Werneck’s survey [36] provides a detailed account of POI recommendation techniques from 2017 to 2019, offering valuable insights into the evolution of methodologies during that time frame, but it lacks coverage of more recent advancements such as the integration of GNNs and LLMs, which bring about not only powerful user modeling capabilities but also intensive computation and scalability issues. To summarize, a noticeable gap in the literature regarding cutting-edge models, architectural evolution, and security considerations is shown in Table I.
TABLE I: In comparing our survey to related surveys in the field, several key distinctions and contributions emerge.
Survey [37] [36] [35] [34] Ours
Taxonomy No Yes Yes Yes Yes
Models Yes Yes Yes Yes Yes
Architecture No No No No Yes
Security No No No No Yes
Year 2016 2020 2022 2023 2024
@@ II Preliminaries
The process of POI recommendation presents a series of challenges centered around identifying suitable upcoming POIs for users based on their historical check-in data and other pertinent information within an LBSN. Consider a group denoted as $U=\left\{u_{1},u_{2},\ldots,u_{N}\right\}$, comprising $N$ users within the LBSN, and a set $P=\left\{p_{1},p_{2},\ldots,p_{M}\right\}$ representing $M$ POIs. Users form connections through a network denoted as $\tilde{U}=\left\{(u_{i},u_{j})|u_{i},u_{j}\in U\right\}$. Each POI $p$ is defined by its geographical coordinates, specifically latitude $x_{p}$ and longitude $y_{p}$, along with a set of attributes $A_{p}$ describing the POI’s semantics. The primary objective is to provide personalized recommendations of relevant POIs based on users’ past interactions and preferences within the LBSN. The key concepts of POI recommendation are articulated as follows:
@@
  • Check-in Event: A check-in event, denoted as $c^{u}_{o}$, involves a user publicly declaring their presence at a specific venue or location at a specific time step $o$. Typically facilitated through mobile applications or social media platforms, this action entails sharing the user’s current whereabouts with their social circle. During a check-in, the user typically selects or searches for a particular place, such as a restaurant, park, or event space, and indicates their current presence at that spot.
  • Next POI Recommendation: Given a user $u$’s check-in list $C^{u}$, the next POI recommendation task revolves around predicting the subsequent POI at time $o_{i+1}$.
@@ III Taxonomy
This paper distinguishes itself from existing surveys on POI recommendation by adopting a holistic approach that emphasizes three critical and interrelated aspects: models, architectures, and security. This tripartite framework, as illustrated in Fig. 2, serves as a comprehensive taxonomy for systematically analyzing and comparing various POI recommendation systems. To further contextualize the research landscape, we present the distribution of existing studies across these three categories in Fig. 1 (a). The figure underscores a clear trend: methods for improving POI recommendation models dominate the field, comprising over 75% of the total research efforts. This focus reflects the ongoing challenge of enhancing accuracy, personalization, and context-awareness in POI recommendations. On the other hand, studies addressing security issues in POI recommendations remain underrepresented, making up approximately 8% of the existing research. This highlights a significant gap in the literature, given the increasing importance of safeguarding user privacy, defending against adversarial attacks, and ensuring data integrity within location-based services. In contrast, research on architectures, which bridges the gap between models and real-world deployment, occupies the remaining portion, indicating its crucial yet often overlooked role in ensuring the scalability, efficiency, and responsiveness of POI recommendation. In the rest of this section, the introduction of the taxonomy will unfold along these three dimensions.
@@ III-A Models
Building on previous POI recommendation surveys [38, 39], which primarily reviewed studies based on traditional non-deep learning and early deep learning methods, there remains a need to address the advances brought by more recent techniques. With the emergence of novel approaches such as diffusion models, self-supervised learning, and large language models, a growing body of research has evolved that is not yet covered in existing surveys. To fill this gap, we provide an extensive review of models spanning from traditional methods to cutting-edge techniques, as illustrated in Fig. 2. To further clarify the progression of these methodologies, we present a visual depiction of their structural evolution in Fig. 1 (b), where the labels are arranged chronologically from bottom to top. This illustration highlights how the research landscape has shifted over time, with an increasing focus on more sophisticated and advanced models. Detailed explanations of each method, along with the corresponding preliminaries, are provided in Section IV.
@@ III-A4 Generative Model-Based
LLMs, renowned for their ability to comprehend and generate human-like text, have emerged as a transformative force in recommendation systems, including POI recommendations [46, 47]. Traditionally, recommendation models relied on structured data like user-item interactions, but LLMs introduce a paradigm shift by incorporating rich, unstructured textual data, including user reviews, POI descriptions, and contextual insights such as time, location, and trends. This allows LLMs to exceed traditional recommendation approaches constrained by interaction data. By processing this diverse information, LLMs can generate more personalized, context-aware, and nuanced recommendations. Furthermore, LLMs enhance user experience by engaging users in natural dialogue, making interactions more intuitive. The POI recommendation generation process with LLMs can be formalized as follows:
@@
DMs, which learn to gradually add noise to data and reverse the process to generate new samples, can be applied to POI recommendation by generating plausible user-POI interaction patterns. They can create new user-POI interaction data, effectively filling in missing entries in the interaction matrix. This addresses data sparsity and improves recommendation accuracy. This research line can be formulated as follows:
@@
VAEs [48, 49] are generative models that learn to represent data in a compressed latent space. They consist of an encoder mapping input data to a distribution in this latent space and a decoder reconstructing data from samples drawn from that distribution. This allows VAEs to generate new instances by sampling from the learned latent space. They enhance POI recommendation systems by capturing user preferences and location characteristics in a compressed space. By encoding user interaction data with various POIs, VAEs capture underlying patterns and preferences, generating personalized recommendations and improving user experience in location-based services.
@@ III-B1 Centralized-Based
III-B2 Decentralized-Based
Decentralized-based structures shift the processing burden from centralized servers to users’ individual devices (e.g., smartphones) [54]. In this approach, recommendation models are deployed locally on the user’s device, where data processing and model inference occur without the need to continuously communicate with a central server. This architecture offers several advantages for POI recommendation, such as enhanced privacy, since sensitive user data remains on the device, and reduced latency, as recommendations can be generated in real-time without relying on server-side processing. Decentralized approaches are particularly suitable for mobile or offline scenarios, where data privacy is critical, or network connectivity is limited. However, the performance of decentralized models may be constrained by the device’s computational resources, and updating models across devices can be challenging. Formally, the optimization of a decentralized model can be formulated as $\theta^{*}=\arg\min_{\theta}\mathcal{L}(\theta;\mathcal{D}_{u})$, where $\theta$ represents the model parameters specific to the device, and $\mathcal{L}(\theta;\mathcal{D}_{u})$ is the loss function computed over the local dataset $\mathcal{D}_{u}$ belonging to the user. By leveraging local computation, decentralized architectures provide a privacy-preserving solution that enables real-time, personalized recommendations tailored to individual user data.
@@ III-C2 User Privacy Protection
where $\mathbf{Enc}(\cdot)$ denotes the encryption algorithm under a specific key, $\oplus$,$\otimes$ , ⊗ represent addition and multiplication over ciphertext, +,+$\cdot$+ , ⋅ represent addition and multiplication over plaintext, and $x,y$ are the input data. This property allows computations to be performed on encrypted data, producing encrypted results that can be decrypted to obtain the final output. These methods facilitate data analysis without compromising the confidentiality of personal information, allowing systems to generate accurate recommendations while maintaining user trust and privacy.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=center, minimum width=1em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [POI Recommendation,leaf-head, ver [ Models (Sec. IV), leaf-task,text width=7em [ Latent Factor Model-Based, leaf-task10, text width=11.5em [LDA-Based, leaf-task8, text width=7.5em[LCARS [64, 16], JIM [65], Geo-SAGE [66], TRM
 [65], LA-LDA [67], TRM+ [68], SPORE [69], ST-LDA [70],
UCGT [71], LSARS [72], ST-SAGE [73], TPM [74], modelnode-task8,text width=29em]] [MF-Based, leaf-task9, text width=7.5em[GeoMF [75], GeoMF++ [76], LGLMF [77], SSTPMF [78],
SQPMF [79], modelnode-task9,text width=29em]] ] [ Classic NN-Based, leaf-taska, text width=9.5em [LSTM-Based, leaf-task1, text width=9.5em [ST-LSTM [80], LSTM-S [81], PS-LSTM [82], STMLA [83], LSMA
 [84], AT-LSTM [85], ATSD-GRU [86], DLAN [87],modelnode-task1, text width=29em] ] [Transformer-Based, leaf-task2, text width=9.5em [TLR-M [41], GETNext [88], CARAN [89], JANICP [90],
Li et al. [91], STUIC-SAN [92], AMACF [93], CHA [94],
MVSAN [95], HAT [96], STAR-HiT [97], CAFPR [98],
FPG [99], TGAT [100], MobGT [101], POIBERT [102],
AutoMTN [103], CCDSA [104], Liu  et al. [105], TDGCN [106],
TGAT [100], MGAN [107] , STA-TCN [108] , Xia et al. [109],
STTF-Recommender [110] , BayMAN [111], Kumar et al. [112],
ADQ-GNN [22] , GN-GCN [119] , ATST-GGNN [120] , Zhao et al.
 [121] , HS-GAT [122] , PROG-HGNN [28] , CGNN-PRRG [123] ,
Gao et al. [130], Wang et al. [131], SLS-REC [132], CLLP [45], , modelnode-task6, text width=40em] ] [ Generative-Based, leaf-taska, text width=9.5em [ LLM-Based, leaf-task5, text width=9.5em [LLMmove [46], Li et al. [47], modelnode-task5,text width=29em] ] [ DM-Based, leaf-task7, text width=9.5em [Diff-POI [133], DSDRec [134] , Diff-DGMM [135] , DCPR [136], modelnode-task7,text width=29em] ] [VAE-Based, leaf-task4, text width=9.5em [WaPOIR [137], Li et al. [138], modelnode-task4, text width=29em] ] ] ] [ Architectures
(Sec. V), leaf-paradigms,text width=7em [ Centralized, leaf-paradigms, text width=9.5em [ GeoIE [51], SH-CDL [50], GETnext [88], STGN [139], AGRAN [53] , modelnode-paradigms, text width=40em] ] [ Decentralized-Based, leaf-paradigms, text width=9.5em [ MAC [54], DARD [140], LLRec [141], DCLR [27], modelnode-paradigms, text width=40em] ] [ FL-Based, leaf-paradigms, text width=9.5em [ FL&PP-POI [142], Huang et al. [143],CPF-POI [144], SCFL [145], SFL [146], FedPoiRec [29],
PrefFedPOI [147], PREFER [148], Dong et al. [149], DMF [60], NRDL [150] , modelnode-paradigms, text width=40em] ] ] [ Security (Sec. VI), leaf-others,text width=7em [ Data Integrity Threats, leaf-others, text width=9.5em [FedRecAttack [56], Ghost riders [58] , LOKI [57], Yuan et al. [151], PTIA [59] , modelnode-others, text width=40em] ] [ User Privacy Protection, leaf-others, text width=9.5em [DMF [152], Li et al. [62], PriRec [61], Liu et al. [63], Huo et al. [153], FedPoiRec [29], modelnode-others, text width=40em] ] ] ]

### v2

@@ I Introduction
The proliferation of smart devices has fueled the rapid growth of Location-based Social Networks (LBSNs) [1, 2, 3], which allow users to share check-ins, reviews, and personal experiences tied to specific locations. These networks, now with billions of users, generate vast amounts of spatio-temporal data [4, 5, 6, 7, 8, 9, 10, 10, 11], presenting valuable opportunities for personalized Point-of-Interest (POI) recommendations. As a dynamic area in recommendation systems, POI recommendation has gained considerable interest from both users and businesses in recent years. These methods leverage users’ historical check-ins along with multimodal data to suggest personalized destinations [6, 12]. However, the diversity in data size, modality, and user expectations introduces new challenges. These complexities motivate researchers to develop innovative techniques that effectively capture mobility patterns and other relevant features, such as spatial, social, and textual information, to enhance the performance of POI recommendations [13, 14, 11].
@@
While existing surveys [35, 36, 37, 38, 39] have contributed valuable insights into POI recommendation systems, there remains a critical need for a comprehensive review that reflects the rapid developments across POI models, architectures, and security. For instance, while Zhao et al. [35] provided a thorough review of POI recommendation with traditional techniques like matrix factorization, the growing challenges within deep POI recommender systems are not covered. By contrast, Wang et al. [38] provided an overview of various POI recommendation methods in the deep learning era, yet they do not delve deeply into the architectural and security challenges that have emerged with the rise of decentralized systems and practical privacy concerns. Similarly, Islam’s survey [37] emphasizes the impact of deep learning on POI recommendation, but it overlooks key advancements in federated learning and edge computing, which are increasingly shaping the deployment of these systems. Furthermore, Werneck’s survey [36] provides a detailed account of POI recommendation techniques from 2017 to 2019, offering valuable insights into the evolution of methodologies during that time frame, but it lacks coverage of more recent advancements such as the integration of GNNs and LLMs, which bring about not only powerful user modeling capabilities but also intensive computation and scalability issues. Recently, Yin et al. [39] provided a comprehensive survey on on-device recommender systems, covering aspects such as deployment, inference, training, updating, and the security and privacy of these systems. Consequently, their work does not address key aspects of POI recommendation, such as the foundational models used. As shown in Table I, there is a noticeable gap in the literature regarding cutting-edge models, architectural evolution, and security considerations.
TABLE I: Comparison of related surveys.
Survey Year Taxonomy Models Architecture Security
[35] 2016 $\times$ $\checkmark$ $\times$ $\times$
[36] 2020 $\checkmark$ $\checkmark$ $\times$ $\times$
[37] 2022 $\checkmark$ $\checkmark$ $\times$ $\times$
[38] 2023 $\checkmark$ $\checkmark$ $\times$ $\times$
[39] 2024 $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Ours 2024 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ II Preliminaries
The process of POI recommendation presents a series of challenges centered around identifying suitable upcoming POIs for users based on their historical check-in data and other pertinent information within an LBSN. Consider a group denoted as $U=\left\{u_{1},u_{2},\ldots,u_{N}\right\}$, comprising $N$ users within the LBSN, and a set $P=\left\{p_{1},p_{2},\ldots,p_{M}\right\}$ representing $M$ POIs. Users form connections through a network denoted as $\tilde{U}=\left\{(u_{i},u_{j})|u_{i},u_{j}\in U\right\}$. Each POI $p$ is defined by its geographical coordinates, specifically latitude $x_{p}$ and longitude $y_{p}$, along with a set of attributes $A_{p}$ describing the POI’s semantics. The primary objective is to provide personalized recommendations of relevant POIs based on users’ past interactions and preferences within the LBSN. Table II provides a concise description of the primary notations used in this paper.
TABLE II: Notations used throughout this paper.
Notation Description
$u_{i}$ User $i$ where $i\in\{1,2,...,N\}$
$U$ Set of $N$ users within the LBSN, where $U=\left\{u_{1},u_{2},\ldots,u_{N}\right\}$.
$p_{i}$ POI, a specific location recognized for its significance or interest.
$P$ Set of $M$ POIs, where $P=\{p_{1},p_{2},...,p_{N}\}$.
$\tilde{U}$ Network of user connections within the LBSN, where $\tilde{U}=\left\{(u_{i},u_{j})|u_{i},u_{j}\in U\right\}$.
$x_{p}$ Latitude of POI $p$.
$y_{p}$ Longitude of POI $p$.
$A_{p}$ Set of attributes describing the semantics of POI $p$.
$c^{u}_{o}$ Check-in event where user $u$ declares presence at a location at time $o$.
$C^{u}$ List of check-in events for user $u$, where $C^{u}=\left\{c^{u}_{1},c^{u}_{2},\dots,c^{u}_{k}\right\}$.
The key concepts of POI recommendation are as follows:
@@
  • Check-in Event: A check-in event, denoted as $c^{u}_{o}$, involves a user publicly declaring their presence at a venue or location at a specific time step $o$. Typically facilitated through mobile apps or social media, this action shares the user’s whereabouts with their social circle. During a check-in, the user selects or searches for a place, such as a restaurant, park, or event space, to indicate their presence.
  • POI Recommendation: Given a user $u$’s check-in list $C^{u}=\left\{c^{u}_{1},c^{u}_{2},\dots,c^{u}_{k}\right\}$, where each $c^{u}_{i}$ represents a check-in event at time $o_{i}$, the POI recommendation task aims to suggest a set of POIs to the user based on their historical check-ins and preferences. The goal is to suggest POIs that the user is likely to visit in the future, leveraging their interaction patterns and preferences within the LBSN.
  • Next POI Recommendation: Given a user $u$’s check-in list $C^{u}$, the next POI recommendation task revolves around predicting the subsequent POI at time $o_{i+1}$. It specifically refers to recommending the immediate next POI the user is likely to visit based on their current trajectory and patterns.
  • Spatial Item Recommendation: Given a user’s historical interaction data at various spatial locations, the spatial item recommendation task aims to suggest items or products that are contextually relevant to the user based on their location and spatial preferences. The goal is to recommend items that the user is likely to engage with, considering both the spatial distribution of items and the user’s past interactions within these geographical regions.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=center, minimum width=1em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [POI Recommendation,leaf-head, ver [ Models (Sec. IV), leaf-task,text width=8em [ Latent Factor Model-Based, leaf-task10, text width=11.5em [LDA-Based, leaf-task8, text width=7.5em[LCARS [40, 17], JIM [41], Geo-SAGE [42], TRM [41],
LA-LDA [43], TRM+ [44], SPORE [45], ST-LDA [46], UCGT [47],
LSARS [48], ST-SAGE [49], TPM [50], modelnode-task8,text width=29em]] [MF-Based, leaf-task9, text width=7.5em[GeoMF [51], GeoMF++ [52], LGLMF [53], SSTPMF [54],
SQPMF [55], modelnode-task9,text width=29em]] ] [ Classic NN-Based, leaf-taska, text width=9.5em [LSTM-Based, leaf-task1, text width=9.5em [ST-LSTM [56], LSTM-S [57], PS-LSTM [58], STMLA [59],
LSMA [60], AT-LSTM [61], ATSD-GRU [62], DLAN [63],modelnode-task1, text width=29em] ] [Transformer-Based, leaf-task2, text width=9.5em [TLR-M [64], GETNext [65], CARAN [66], JANICP [67],
Li et al. [68], STUIC-SAN [69], AMACF [70], CHA [71],
MVSAN [72], HAT [73], STAR-HiT [74], CAFPR [75],
FPG [76], TGAT [77], MobGT [78], POIBERT [79],
AutoMTN [80], CCDSA [81], Liu  et al. [82], TDGCN [83],
TGAT [77], MGAN [84] , STA-TCN [85] , Xia et al. [86],
STTF-Recommender [87] , BayMAN [88], Kumar et al. [89],
ADQ-GNN [23] , GN-GCN [96] , ATST-GGNN [97] , Zhao et al.
 [98] , HS-GAT [99] , PROG-HGNN [29] , CGNN-PRRG [100] ,
Gao et al. [108], Wang et al. [109], SLS-REC [110], CLLP [111], , modelnode-task6, text width=40em] ] [ Generative, leaf-taska, text width=9.5em [ LLM-Based, leaf-task5, text width=9.5em [LLMmove [112], Li et al. [113], modelnode-task5,text width=29em] ] [ DM-Based, leaf-task7, text width=9.5em [Diff-POI [114], DSDRec [115] , Diff-DGMM [116] , DCPR [117], modelnode-task7,text width=29em] ] [VAE-Based, leaf-task4, text width=9.5em [WaPOIR [118], Li et al. [119], modelnode-task4, text width=29em] ] ] ] [ Architectures
(Sec. V), leaf-paradigms,text width=8em [ Centralized, leaf-paradigms, text width=9.5em [ GeoIE [120], SH-CDL [121], GETnext [65], STGN [122], AGRAN [123] , modelnode-paradigms, text width=40em] ] [ Decentralized, leaf-paradigms, text width=9.5em [ MAC [124], DARD [125], LLRec [126], DCLR [28], modelnode-paradigms, text width=40em] ] [ FL-Based, leaf-paradigms, text width=9.5em [ FL&PP-POI [127], Huang et al. [128],CPF-POI [129], SCFL [130], SFL [131], FedPoiRec [30],
PrefFedPOI [132], PREFER [133], Dong et al. [134], DMF [135], NRDL [136] , modelnode-paradigms, text width=40em] ] ] [ Security (Sec. VI), leaf-others,text width=8em [ System Vulnerability
Analysis, leaf-others, text width=9.5em [FedRecAttack [137], Ghost riders [138] , LOKI [139], Yuan et al. [140], PTIA [141] , modelnode-others, text width=40em] ] [ User Privacy Protection, leaf-others, text width=9.5em [DMF [135], Li et al. [142], PriRec [143], Liu et al. [144], Huo et al. [145], FedPoiRec [30], modelnode-others, text width=40em] ] ] ]
@@ III Taxonomy
This paper distinguishes itself from existing surveys on POI recommendation by adopting a holistic approach that focuses on three critical, interrelated aspects: models, architectures, and security. This tripartite framework serves as a comprehensive taxonomy for analyzing and comparing various POI recommendation systems.
To contextualize the research landscape, Fig. 1 illustrates the framework, while Fig. 2 (a) presents the distribution of existing studies across these three categories. Research on POI recommendation models dominates the field, accounting for over 75% of studies, reflecting ongoing challenges in improving accuracy, personalization, and context-awareness. In contrast, security-related research remains limited, comprising around 8% of the literature, underscoring gaps in addressing user privacy, adversarial threats, and data integrity in location-based services. Research on architectures, which ensures scalability and efficiency in real-world deployments, occupies the remaining portion, emphasizing its critical yet often overlooked role. To further illustrate the evolution of methodologies in POI recommendation, Fig. 2 (b) provides a structured timeline, where key techniques are arranged chronologically from bottom to top. This visualization highlights the shift from traditional models such as LDA and MF to more advanced techniques, including LSTMs, GNNs, transformers, and large language models (LLMs). Additionally, it categorizes architectures into centralized, decentralized, and federated-learning-based approaches, while security-related concerns are mapped to areas such as user privacy protection and system vulnerability analysis. This progression underscores the increasing sophistication and diversification of POI recommendation research over time.
@@ III-A Models
Building on previous POI recommendation surveys [146, 147], which primarily reviewed studies based on traditional non-deep learning and early deep learning methods, we address the advancements brought by more recent techniques. With the rise of novel approaches such as diffusion models, self-supervised learning, and large language models, a growing body of research remains uncovered in existing surveys. To bridge this gap, we provide an extensive review of models from traditional methods to cutting-edge techniques. Detailed explanations of each method, along with the corresponding preliminaries, are provided in Section IV.
@@ III-A4 Generative
Large language models, renowned for their ability to comprehend and generate human-like text, have emerged as a transformative force in recommendation systems, including POI recommendations [112, 113]. Traditionally, recommendation models relied on structured data like user-item interactions, but LLMs introduce a paradigm shift by incorporating rich, unstructured textual data, including user reviews, POI descriptions, and contextual insights such as time, location, and trends. This allows LLMs to exceed traditional recommendation approaches constrained by interaction data. By processing this diverse information, LLMs can generate more personalized, context-aware, and nuanced recommendations. Furthermore, LLMs enhance user experience by engaging users in natural dialogue, making interactions more intuitive. The POI recommendation generation process with LLMs can be formalized as follows:
@@
Diffusion models, which learn to gradually add noise to data and reverse the process to generate new samples, can be applied to POI recommendation by generating plausible user-POI interaction patterns. They can create new user-POI interaction data, effectively filling in missing entries in the interaction matrix. This addresses data sparsity and improves recommendation accuracy. This research line can be formulated as follows:
@@
Variational autoencoders [151, 152] are generative models that learn to represent data in a compressed latent space. They consist of an encoder mapping input data to a distribution in this latent space and a decoder reconstructing data from samples drawn from that distribution. This allows VAEs to generate new instances by sampling from the learned latent space. They enhance POI recommendation systems by capturing user preferences and location characteristics in a compressed space. By encoding user interaction data with various POIs, VAEs capture underlying patterns and preferences, generating personalized recommendations and improving user experience in location-based services.
@@ III-B1 Centralized-Based
III-B2 Decentralized
Decentralized architectures shift the processing burden from centralized servers to users’ individual devices (e.g., smartphones) [124]. In this approach, recommendation models are deployed locally on the user’s device, where data processing and model inference occur without the need to continuously communicate with a central server. This architecture offers several advantages for POI recommendation, such as enhanced privacy, since sensitive user data remains on the device, and reduced latency, as recommendations can be generated in real-time without relying on server-side processing. Decentralized approaches are particularly suitable for mobile or offline scenarios, where data privacy is critical, or network connectivity is limited. However, the performance of decentralized models may be constrained by the device’s computational resources, and updating models across devices can be challenging. Formally, the optimization of a decentralized model can be formulated as $\theta^{*}=\arg\min_{\theta}\mathcal{L}(\theta;\mathcal{D}_{u})$, where $\theta$ represents the model parameters specific to the device, and $\mathcal{L}(\theta;\mathcal{D}_{u})$ is the loss function computed over the local dataset $\mathcal{D}_{u}$ belonging to the user. By leveraging local computation, decentralized architectures provide a privacy-preserving solution that enables real-time, personalized recommendations tailored to individual user data.
@@ III-C2 User Privacy Protection
where $\mathbf{Enc}(\cdot)$ denotes the encryption algorithm under a specific key, $\oplus$ and $\otimes$ represent addition and multiplication over ciphertext, $+$ and $\cdot$ represent addition and multiplication over plaintext, and $x,y$ are the input data. This property allows computations to be performed on encrypted data, producing encrypted results that can be decrypted to obtain the final output. These methods facilitate data analysis without compromising the confidentiality of personal information, allowing systems to generate accurate recommendations while maintaining user trust and privacy.